An electron enters between two horizontal plates seprated by 2mm and having a potential difference of 1000v . The force on electron is 

Grade:12

1 Answers

Pratham Ashish
14 years ago

hi,

let electric field strength be E

 V = E.d

E = 1000/ 2* 10 ^-3

   = 500* 10 ^3 V/m

 force on electron,

  F = qE

     =1.6 X 10^-19  X  500* 10 ^3

     = 8.0 * 10^-14 N